# The disk precession in a Be star-magnetar binary and its application to the rotation measure of FRB 20201124A

### v1

@@ 5.2 Conclusions
In this work, we demonstrated that a Be star-magnetar binary system with a decretion stellar disk is suitable for explaining the peculiar RM variations observed by FAST from MJD 59307 to MJD 59658.
@@
The RM in the first epoch exhibits a pattern of rising, a deeper drop, rising again, and then entering a steady phase. Such a pattern strongly supports that FRB 20201124A originates from a binary system consisting of a magnetar and a massive star with a decretion disk. In the later two observation epochs, the RM fluctuations decrease significantly compared with $\Delta\rm{RM_{E1}}$. The precession of the Be star along with its decretion disk is commonly used to explain the observed superorbital brightness oscillations of Be/X-ray binary system.
@@
We propose that the long-term decrease in the observed RM variations can be interpreted by the superorbital modulation from a precessing stellar disk in a Be star-magnetar binary. By fitting the data, we obtainthe bestt parameters that fitthe RMM curves. The best parameters are presented in Section 4.
@@
We appreciate Dr. A. M. Chen for useful discussion of the precession. This work is supported by the National Natural Science Foundation of China under grants 12041306, 12473012, 12533005, 12173014 12494575, 12273009 and 12573044, and the National Key R&D Program of China (Nos. 2023YFC2205901, 2020YFC2201400, SQ2023YFC220007). W.H.Lei. acknowledges support from science research grants from the China Manned Space Project with NO.CMS-CSST-2021-B11.
Table 2: Parameters of fitting curves for the three FRB sources in Figure 3
Parameters FRB 20180301A FRB 20190520A FRB 20191106C
$M_{\star}\,(M_{\odot})$ 8 8 8
$R_{\star}\,(R_{\odot})$ 5 5 5
$M_{\rm{NS}}\,(M_{\odot})$ 1.6 1.6 1.6
$\varepsilon$ 0.5 0.5 0.7
$P_{\rm{orb}}\,\rm{(day)}$ 2500 800 4500
$\rho_{0}\,\rm{(g\,cm^{-3})}$ $9.0\times 10^{-14}$ $6.2\times 10^{-12}$ $4.0\times 10^{-13}$
$\beta$ 2.5 3.0 3.0
$\varphi\,\rm{(rad)}$ -0.785 -0.785 -0.780
$\theta\,\rm{(rad)}$ 0.1 0.01 0.05
$\omega\,\rm{(rad\,day^{-1})}$ $1.26\times 10^{-3}$ $3.93\times 10^{-3}$ $6.98\times 10^{-4}$
$\lambda_{1}\,\rm{(rad)}$ -0.61 -0.40 -0.35
$\lambda_{2}\,\rm{(rad)}$ 3 2.98 2.6

### v2

@@ V.2 Conclusions
In this work, we demonstrate that a Be star-magnetar binary system with a decretion stellar disk is suitable for explaining the peculiar RM variations observed by FAST from MJD 59307 to MJD 59658.
@@
The RM in the first epoch exhibits a pattern of rising, a deeper drop, rising again, and then entering a steady phase. Such a pattern strongly supports that FRB 20201124A originates from a binary system consisting of a magnetar and a massive star with a decretion disk. In the subsequent two observation epochs, the RM fluctuations decrease significantly compared with $\Delta\rm{RM_{E1}}$. The precession of the Be star along with its decretion disk is commonly used to explain the observed superorbital brightness oscillations of Be/X-ray binary system.
@@
We propose that the long-term decrease in the observed RM variations can be interpreted by the superorbital modulation from a precessing stellar disk in a Be star-magnetar binary. By fitting the data, we obtain the best parameters that fit the RM curves. The best parameters are presented in Section IV.
@@
We thank the anonymous referee for detailed and very constructive suggestions that have allowed us to improve our manuscript. We appreciate Dr. A. M. Chen for useful discussion of the precession. This work is supported by the National Natural Science Foundation of China under grants 12041306, 12473012, 12533005, 12173014 12494575, 12273009 and 12573044, and the National Key R&D Program of China (Nos. 2023YFC2205901, 2020YFC2201400, SQ2023YFC220007). W.H.Lei. acknowledges support from science research grants from the China Manned Space Project with NO.CMS-CSST-2021-B11.